OpenStudy (anonymous):

Solve the logarithmic equation. log2(x-1) - log2(5x + 1) = -3 log2[(x-1)(5x + 1)] = -3 (x-1)(5x + 1) = 2^-3 5x^2 - 5x - 1 = 1/8 5x^2 + 4x - 7/8 = 0 This class is killing me.

OpenStudy (anonymous):

I was trying to follow along with my book. But of course the only give easy examples.

OpenStudy (anonymous):

is it the log to base 2?

OpenStudy (anonymous):

log2+log(x-1)-log2-log(5x+1)=-3 or, log(x-1)-log(5x+1)=-3 or, log((x-1)/(5x+1))=-3 or, (x-1)/(5x+1)=2pow(-3)=1/8 now try to find the value of x

OpenStudy (anonymous):

yes.